OpenStudy (anonymous):

Write the equation of the line that passes through (2, 2) and (6, 3) in standard form. 4x + 4y = 6 x + 4y = 12 4x + 4y = –12 x – 4y = –6 I have this so far I think it might be wrong. 3-2 =1 6-2 =4 Y-3 = 1/4(x-6) Y-3 = 1/4x + 1.5

hartnn (hartnn):

just a sign mistake : Y-3 = 1/4(x-6) Y-3 = 1/4x - 1.5 rest everything is correct.

hartnn (hartnn):

then multiply 4 on both sides.

OpenStudy (anonymous):

y-12=4x-1.5 -4x + y-12 = -1.5 -4x + y= 10.5 Awnser: 4x – y = -10.5

hartnn (hartnn):

when u multiplied 4, 4*(1/4)x = x and not 4x

hartnn (hartnn):

and 4*y=4y

OpenStudy (anonymous):

let me try it again

OpenStudy (anonymous):

4y-3 = x -1.5 -x + 4y – 3 = -1.5 -x + 4y = 2.5 X – 4y = -2.5

OpenStudy (anonymous):

But its still not an awnser

hartnn (hartnn):

wait i will write the steps.... 4(y-3)= 4/(1/4)x -4*1.5 4y-12 = x - 6 x-4y = -12+6 x-4y = -6 ask if there is doubt in any step..........

OpenStudy (anonymous):

OK so i had to multiply the y And the 3. OK now I see where I was going wrong thank you for the help and being patient with me.

hartnn (hartnn):

welocme :)